Please help!
Find the value of the function.
Given f(x)=1/4x+5, find f(-3)

Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(69443)

Find the value of the function.
Given f(x)=1/4x+5, find f(-3)
-----------
Sub -3 for x.

Answer by stanbon(75887)

Find the value of the function.
Given f(x)=1/4x+5, find f(-3)
------
f(-3) = (1/4)(-3) + 5
---
f(-3) = -3/4 + (20/4)
---
f(-3) = 17/4
================
Cheers,
Stan H.
================